### v1

# DM-MPPI: Datamodel for Efficient and Safe Model Path Integral Control

Abstract
We extend the Datamodels framework from machine learning to Model Predictive Path Integral (MPPI) control. While datamodels estimate sample influence via regression on fixed datasets, our approach learns to predict influence from sample cost features—enabling real-time estimation for new samples without online regression. The predictor is trained offline on influence coefficients computed via the datamodel framework across diverse MPPI instances, then deployed for efficient sample pruning and adaptive constraint handling. A single learned model addresses both efficiency and safety: low-influence samples are pruned to reduce computation, while monitoring constraint-violation influence enables adaptive penalty tuning. Experiments on path tracking with obstacle avoidance demonstrate $5\times$ sample reduction while maintaining performance and improving constraint satisfaction.
I Introduction
Model Predictive Path Integral (MPPI) control has emerged as a powerful framework for sampling-based optimal control in nonlinear systems [1, 2]. By generating trajectory samples and computing importance-weighted averages, MPPI avoids linearization while achieving real-time performance through parallelization [3]. However, two fundamental challenges remain. First, sample efficiency: many samples may contribute little to the final control, wasting computation. Second, constraint handling: soft constraint penalties are the standard approach [18], but understanding how constraint-violating samples influence the control output remains opaque.
The Datamodels framework from machine learning [8] offers a principled approach to understanding how individual data points affect model outputs. We adapt this framework to MPPI by learning to predict trajectory cost from sample inclusion. The key insight is simple: by predicting total cost, which includes constraint violation penalties, a single datamodel captures both task performance and constraint-relevant influence. The relative importance of constraints can be adjusted through the penalty coefficient without requiring separate models.
I-A Related Work
MPPI and Path Integral Control. Path integral methods derive optimal control laws through exponential weighting of sampled trajectories [4]. Williams et al. [1, 2] developed MPPI for real-time control with GPU parallelization, demonstrating aggressive autonomous driving [3]. Recent variants address robustness [5], smoothness [6], and sample efficiency through unscented guidance [7]. Our work complements these advances by providing interpretable influence analysis.
Data Attribution in Machine Learning. Understanding how training data affects model predictions is a fundamental problem in ML. Influence functions [9] adapt classical robust statistics to deep learning, while Data Shapley [10] applies game-theoretic valuation. TracIn [11] provides efficient gradient-based attribution. The Datamodels framework [8] learns linear predictors of model outputs from training set composition, achieving high accuracy on image classification. TRAK [12] scales attribution to large models via random projections. We are the first to apply datamodels to sampling-based control.
Constraint Handling in MPPI. Standard MPPI handles constraints through soft penalties in the cost function, but this approach cannot guarantee constraint satisfaction and requires careful tuning of penalty weights. To address this limitation, several MPPI-specific methods have been developed. Risk-Aware MPPI (RA-MPPI) [13] replaces the expected cost with Conditional Value-at-Risk (CVaR), steering optimization toward the worst-case tail of the rollout distribution to improve safety in autonomous racing. Shield-MPPI [14] integrates Control Barrier Functions (CBFs) into MPPI through a dual-layer structure that combines cost penalization with a gradient-based local repair step, significantly reducing constraint violations while maintaining computational efficiency on CPUs. For hard constraint satisfaction, Constrained Covariance Steering MPPI (CCS-MPPI) [15] enforces state and control constraints through a tube-based formulation, though it operates probabilistically. Gandhi et al. [22] develop safe importance sampling that biases trajectory generation toward constraint-satisfying regions while preserving theoretical guarantees. Our datamodel-based approach complements these methods by diagnosing how constraint-violating samples influence the control output, enabling principled tuning of penalty parameters rather than enforcing constraints directly.
Learning for Control. Deep learning increasingly intersects with control theory [23, 24]. Data-driven methods can directly synthesize controllers from data [25]. Neural certificates [26] learn Lyapunov and barrier functions alongside policies. Interpretability of learned policies is an emerging concern [27]. Our work contributes to this intersection by bringing data attribution tools to sampling-based control.
I-B Contributions
This paper extends the Datamodels framework to MPPI by learning to predict sample influence from cost features rather than computing influence for fixed datasets. Our key contribution is an offline-online architecture: ground-truth influence coefficients are computed via datamodel regression during training, then predicted via a neural network at runtime—enabling real-time estimation without online regression. A single learned predictor addresses both sample efficiency through pruning and constraint handling through adaptive penalty tuning.
The remainder of this paper is organized as follows. Section II provides background on MPPI control and the Datamodels framework. Section III presents the proposed MPPI-Datamodel framework, detailing the offline training procedure for learning the influence predictor and the online inference phase for sample pruning and adaptive constraint handling. Section IV evaluates the framework on a path tracking task with obstacle avoidance, demonstrating sample efficiency gains and improved safety margins. Section V discusses limitations and directions for future work, and Section VI concludes the paper.
II Preliminaries
II-A Model Predictive Path Integral Control
Consider a nonlinear stochastic system with control-affine dynamics
$dx=f(x)dt+G(x)(u+\epsilon)dt$ (1)
where $x\in\mathbb{R}^{n}$ is the state, $u\in\mathbb{R}^{m}$ is the control input, $f:\mathbb{R}^{n}\to\mathbb{R}^{n}$ is the drift dynamics, $G:\mathbb{R}^{n}\to\mathbb{R}^{n\times m}$ is the control matrix, and $\epsilon\sim\mathcal{N}(0,\Sigma)$ represents control noise.
The total trajectory cost consists of multiple components
$S(\tau)=S_{\text{goal}}(\tau)+S_{\text{control}}(\tau)+\rho\cdot S_{\text{violation}}(\tau)$ (2)
where the goal cost captures task objectives $S_{\text{goal}}(\tau)=\phi(x_{T})+\sum_{t=0}^{T-1}q(x_{t})$, the control cost penalizes actuation $S_{\text{control}}(\tau)=\sum_{t=0}^{T-1}\frac{1}{2}u_{t}^{\top}Ru_{t}$, and the violation cost captures soft constraint penalties $S_{\text{violation}}(\tau)=\sum_{t=0}^{T-1}c(x_{t})$ with $\rho>0$ controlling constraint importance.
Given $K$ sampled trajectories $\{\tau_{k}\}_{k=1}^{K}$, each generated by applying perturbed controls $u_{t}^{k}=\bar{u}_{t}+\epsilon_{t}^{k}$, the optimal control update is
$u_{t}^{*}=\bar{u}_{t}+\sum_{k=1}^{K}w_{k}\epsilon_{t}^{k}$ (3)
where the importance weights follow the Gibbs distribution
$w_{k}=\frac{\exp(-S(\tau_{k})/\lambda)}{\sum_{j=1}^{K}\exp(-S(\tau_{j})/\lambda)}.$ (4)
The temperature $\lambda>0$ governs weight concentration [2].
II-B The Datamodels Framework
The Datamodels framework [8] provides a principled approach to understanding how individual data points influence model outputs. The core idea is to learn a simple predictor that maps training set composition to model behavior, enabling post-hoc attribution of predictions to specific training examples.
For a dataset $\mathcal{S}=\{z_{1},\ldots,z_{n}\}$ and a model output of interest $f(x;S^{\prime})$ trained on subset $S^{\prime}\subseteq\mathcal{S}$, a linear datamodel takes the form
$g_{\theta}(\mathbf{1}_{S^{\prime}})=\theta^{\top}\mathbf{1}_{S^{\prime}}+\theta_{0}$ (5)
where $\mathbf{1}_{S^{\prime}}\in\{0,1\}^{n}$ is a binary vector indicating which data points are included in subset $S^{\prime}$. The coefficient $\theta_{j}$ quantifies the marginal influence of data point $z_{j}$ on the output: positive $\theta_{j}$ indicates that including $z_{j}$ increases the predicted output, while negative $\theta_{j}$ indicates a decreasing effect.
Training proceeds by constructing many random subsets $\{S^{\prime}_{i}\}_{i=1}^{M}$, each formed by including each data point independently with probability $\alpha\in(0,1)$. For each subset, the true model output $f(x;S^{\prime}_{i})$ is computed. The datamodel parameters are then learned via LASSO regression [12]:
$\theta^{*}=\operatorname*{arg\,min}_{\theta}\frac{1}{M}\sum_{i=1}^{M}\left(g_{\theta}(\mathbf{1}_{S^{\prime}_{i}})-f(x;S^{\prime}_{i})\right)^{2}+\mu\|\theta\|_{1}$ (6)
The $\ell_{1}$ penalty encourages sparsity, reflecting the intuition that only a subset of data points meaningfully influence any given prediction.
Unlike influence functions [9], which rely on local approximations around the full training set, datamodels directly learn global input-output relationships across the space of possible training subsets.
II-C Why Linear Models Suit MPPI
The weighted cost can be written as
$\bar{S}(\mathbf{b})=\frac{\sum_{k}b_{k}e^{-S_{k}/\lambda}S_{k}}{\sum_{j}b_{j}e^{-S_{j}/\lambda}}=\frac{N(\mathbf{b})}{D(\mathbf{b})}$ (7)
where both numerator $N(\mathbf{b})$ and denominator $D(\mathbf{b})$ are linear in the inclusion vector $\mathbf{b}$, but their ratio is not. However, when subsets are formed by including each sample independently with probability $\alpha$, the denominator concentrates around its expectation for sufficiently large $\alpha K$:
$D(\mathbf{b})\approx\mathbb{E}[D]=\alpha\sum_{k}e^{-S_{k}/\lambda}.$ (8)
Substituting this approximation yields
$\bar{S}(\mathbf{b})\approx\frac{1}{\mathbb{E}[D]}\sum_{k}b_{k}e^{-S_{k}/\lambda}S_{k}=\sum_{k}\theta_{k}b_{k}$ (9)
where $\theta_{k}\propto e^{-S_{k}/\lambda}S_{k}$. Thus, when subset size is large enough for the denominator to concentrate, the weighted cost becomes approximately linear in $\mathbf{b}$. This approximation degrades when $\alpha K$ is small or when a few samples dominate the weights (low $\lambda$, high cost variance), causing $D(\mathbf{b})$ to fluctuate significantly across subsets. Nonetheless, even when these conditions are not fully met, the linear datamodel can still serve as a practical approximation worth investigating.
III Method: MPPI-Datamodel Framework
[…]
Figure 1: MPPI-Datamodel workflow. Offline phase (green): collect MPPI instances, fit datamodels via LASSO regression, and train influence predictor $h_{\phi}$. Online phase: predict influence using $h_{\phi^{*}}$ with cost features and instance statistics, prune low-influence samples for efficiency, and adapt penalty $\rho$ based on violation influence ratio for safety.
III-A Problem Formulation
Consider an MPPI instance with $K$ sampled trajectories $\mathcal{T}=\{\tau_{1},\ldots,\tau_{K}\}$. Each trajectory $\tau_{k}$ incurs a total cost $C_{k}$ composed of goal cost, control cost, and constraint violation cost:
$C_{k}=C_{\text{goal},k}+C_{\text{ctrl},k}+\rho\cdot C_{\text{viol},k}$ (10)
where $\rho>0$ is the soft constraint penalty coefficient. For a subset $\mathcal{T}^{\prime}\subseteq\mathcal{T}$, the subset-restricted weighted cost is
$\bar{C}(\mathcal{T}^{\prime})=\sum_{k:\tau_{k}\in\mathcal{T}^{\prime}}w_{k}(\mathcal{T}^{\prime})\,C_{k}$ (11)
with renormalized importance weights
$w_{k}(\mathcal{T}^{\prime})=\frac{\exp(-C_{k}/\lambda)}{\sum_{j:\tau_{j}\in\mathcal{T}^{\prime}}\exp(-C_{j}/\lambda)}$ (12)
To capture the relationship between sample inclusion and weighted cost, we define the binary inclusion vector $\mathbf{b}\in\{0,1\}^{K}$ where $b_{k}=1$ if trajectory $\tau_{k}$ is included. The MPPI-Datamodel predicts weighted cost from sample membership via the linear model
$g_{\theta}(\mathbf{b})=\theta^{\top}\mathbf{b}+\theta_{0}\approx\bar{C}(\mathcal{T}^{\prime})$ (13)
where the coefficient $\theta_{k}$ represents the marginal influence of sample $\tau_{k}$ on the weighted cost. Since the MPPI control update $u^{*}=\sum_{k}w_{k}u_{k}$ uses the same importance weights $w_{k}$, samples that significantly affect the weighted cost also significantly affect the control output. This justifies using influence on weighted cost as a proxy for influence on the control decision.
III-B Offline Training Phase
We adopt an offline training paradigm to avoid repeated regression during deployment. The key insight from the datamodeling framework [8] is that by training on many random subsets, we can recover individual sample influences even though we never isolate single samples. The offline procedure consists of three steps.
Step 1: Data Collection. We collect $N$ MPPI instances across representative states sampled from the expected operating distribution. For each instance $i$, we draw $K$ trajectories using the nominal sampling distribution and record the cost components $(C_{\text{goal},k}^{(i)},C_{\text{ctrl},k}^{(i)},C_{\text{viol},k}^{(i)})$ for each trajectory $k$. Given a fixed penalty $\rho_{0}$, we compute total costs $C_{k}^{(i)}=C_{\text{goal},k}^{(i)}+C_{\text{ctrl},k}^{(i)}+\rho_{0}\cdot C_{\text{viol},k}^{(i)}$. To ensure the learned predictor generalizes, instances should cover diverse regions of the state space, including states near constraints.
Step 2: Datamodel Fitting. For each instance $i$, we learn the influence coefficients by regression on random subsets. We generate $M$ random subsets by including each trajectory independently with probability $\alpha\in(0,1)$. For each subset $j$ with inclusion mask $\mathbf{b}_{j}\in\{0,1\}^{K}$, we compute the subset-restricted weighted cost
$\bar{C}_{j}^{(i)}=\frac{\sum_{k=1}^{K}b_{j,k}\,w_{k}^{(i)}\,C_{k}^{(i)}}{\sum_{k=1}^{K}b_{j,k}\,w_{k}^{(i)}}$ (14)
where $w_{k}^{(i)}=\exp(-C_{k}^{(i)}/\lambda)$ are the unnormalized importance weights. The influence coefficients are obtained by solving the LASSO regression problem
$\theta^{(i)*}=\operatorname*{arg\,min}_{\theta,\theta_{0}}\frac{1}{M}\sum_{j=1}^{M}\left(\theta^{\top}\mathbf{b}_{j}+\theta_{0}-\bar{C}_{j}^{(i)}\right)^{2}+\mu\|\theta\|_{1}$ (15)
where $\theta_{0}$ is an intercept term excluded from regularization and the $\ell_{1}$ penalty encourages sparsity.
The key to recovering individual influences is the randomness of the subsets. For any two trajectories $k\neq k^{\prime}$, their inclusion events are independent across subsets:
$\mathbb{E}[b_{j,k}\cdot b_{j,k^{\prime}}]=\mathbb{E}[b_{j,k}]\cdot\mathbb{E}[b_{j,k^{\prime}}]=\alpha^{2}$ (16)
This independence decorrelates the contributions of different samples, allowing the regression to disentangle individual effects from aggregate subset effects. With sufficient subsets $M$, the LASSO solution $\theta_{k}^{(i)*}$ accurately estimates the marginal influence of each sample $k$.
Step 3: Influence Predictor Training. We train a predictor $h_{\phi}$ that generalizes influence estimation across instances. A critical observation is that the influence coefficient $\theta_{k}^{(i)*}$ depends not only on sample $k$’s own cost, but also on the cost distribution of other samples in instance $i$. Intuitively, a trajectory with cost $C_{k}=10$ has high influence if all other trajectories have costs above $100$, but low influence if other trajectories have similar costs around $10$.
To capture this context dependence, we augment the input features with summary statistics:
$\hat{\theta}_{k}=h_{\phi}\left(C_{k},C_{\text{viol},k},\bar{C}^{(i)},\sigma_{C}^{(i)}\right)$ (17)
where $\bar{C}^{(i)}=\frac{1}{K}\sum_{k=1}^{K}C_{k}^{(i)}$ is the mean cost and $\sigma_{C}^{(i)}=\text{std}(\{C_{k}^{(i)}\}_{k=1}^{K})$ is the standard deviation. These statistics provide context about where sample $k$ sits relative to other samples in the instance.
The predictor parameters are learned by minimizing prediction error across all collected instances:
$\phi^{*}=\operatorname*{arg\,min}_{\phi}\sum_{i=1}^{N}\sum_{k=1}^{K}\left(h_{\phi}\left(C_{k}^{(i)},C_{\text{viol},k}^{(i)},\bar{C}^{(i)},\sigma_{C}^{(i)}\right)-\theta_{k}^{(i)*}\right)^{2}$ (18)
In practice, $h_{\phi}$ can be a small neural network (e.g., two hidden layers with 64 units) or even a linear model if the relationship is sufficiently simple. Algorithm 1 summarizes the complete offline training procedure.
Algorithm 1 Offline: Influence Predictor Training
0: Number of instances NN, samples per instance $K$, subsets per instance $M$, inclusion probability $\alpha$, regularization $\mu$, initial penalty $\rho_{0}$
0: Trained predictor $h_{\phi^{*}}$
1: Initialize dataset $\mathcal{D}\leftarrow\emptyset$
2:for $i=1$ to NN do
3:  Sample state $x_{0}^{(i)}$ from operating distribution
4:  Collect $K$ trajectories, compute cost components $\{C_{\text{goal},k}^{(i)},C_{\text{ctrl},k}^{(i)},C_{\text{viol},k}^{(i)}\}_{k=1}^{K}$
5:  Compute total costs: $C_{k}^{(i)}\leftarrow C_{\text{goal},k}^{(i)}+C_{\text{ctrl},k}^{(i)}+\rho_{0}\cdot C_{\text{viol},k}^{(i)}$
6:  Compute importance weights: $w_{k}^{(i)}\leftarrow\exp(-C_{k}^{(i)}/\lambda)$
7:  Compute instance statistics: $\bar{C}^{(i)}\leftarrow\frac{1}{K}\sum_{k}C_{k}^{(i)}$,  $\sigma_{C}^{(i)}\leftarrow\text{std}(\{C_{k}^{(i)}\})$
8:  for $j=1$ to $M$ do
9:   Sample inclusion mask: $b_{j,k}\sim\text{Bernoulli}(\alpha)$ for all $k$
10:   Compute weighted cost: $\bar{C}_{j}^{(i)}\leftarrow\frac{\sum_{k}b_{j,k}\,w_{k}^{(i)}\,C_{k}^{(i)}}{\sum_{k}b_{j,k}\,w_{k}^{(i)}}$
11:  end for
12:  Solve LASSO (15) to obtain $\theta^{(i)*}\in\mathbb{R}^{K}$
13:  for $k=1$ to $K$ do
14:   $\mathcal{D}\leftarrow\mathcal{D}\cup\{(C_{k}^{(i)},C_{\text{viol},k}^{(i)},\bar{C}^{(i)},\sigma_{C}^{(i)},\theta_{k}^{(i)*})\}$
15:  end for
16:end for
17: Train $h_{\phi^{*}}$ on $\mathcal{D}$ by minimizing (18)
18:return $h_{\phi^{*}}$
III-C Online Inference Phase
At runtime, the learned predictor enables efficient influence estimation without solving any regression problems. The predictor uses the decomposed costs $(C_{k},C_{\text{viol},k})$ along with instance statistics as input features. Importantly, the actual MPPI control computation still operates on the total cost with the current penalty $\rho$, which can be adjusted online without retraining the predictor.
Given a new set of $K$ sampled trajectories, we first compute the cost components $(C_{\text{goal},k},C_{\text{ctrl},k},C_{\text{viol},k})$ and total costs $C_{k}=C_{\text{goal},k}+C_{\text{ctrl},k}+\rho\cdot C_{\text{viol},k}$. We then compute the instance statistics
$\bar{C}=\frac{1}{K}\sum_{k=1}^{K}C_{k},\qquad\sigma_{C}=\text{std}(\{C_{k}\}_{k=1}^{K})$ (19)
and obtain the predicted influence coefficients as
$\hat{\theta}_{k}=h_{\phi^{*}}(C_{k},C_{\text{viol},k},\bar{C},\sigma_{C})$ (20)
These predictions identify which samples contribute meaningfully to the control output.
Sample Pruning for Efficiency. We retain only those samples whose predicted influence magnitude exceeds a threshold $\tau$:
$\mathcal{K}_{\text{keep}}=\{k:|\hat{\theta}_{k}|\geq\tau\}$ (21)
This pruning step reduces computational cost while preserving the samples that matter most. The MPPI control update is then computed using only the retained samples:
$u^{*}=\sum_{k\in\mathcal{K}_{\text{keep}}}\tilde{w}_{k}\,u_{k},\qquad\tilde{w}_{k}=\frac{\exp(-C_{k}/\lambda)}{\sum_{j\in\mathcal{K}_{\text{keep}}}\exp(-C_{j}/\lambda)}$ (22)
where the importance weights are renormalized over $\mathcal{K}_{\text{keep}}$.
Adaptive Constraint Penalty for Safety. To adapt the constraint penalty online, we monitor the influence contributed by constraint-violating samples. We define the violation influence ratio as
$r_{\text{viol}}=\frac{\sum_{k\in\mathcal{K}_{\text{viol}}}|\hat{\theta}_{k}|}{\sum_{k=1}^{K}|\hat{\theta}_{k}|}$ (23)
where $\mathcal{K}_{\text{viol}}=\{k:C_{\text{viol},k}>0\}$ denotes the set of constraint-violating samples. This ratio quantifies how much of the total influence comes from samples that violate constraints.
When $r_{\text{viol}}$ is high, constraint-violating trajectories disproportionately affect the control output, suggesting the current penalty $\rho$ is insufficient to discourage unsafe behaviors. We update the penalty coefficient according to
$\rho_{t+1}=\rho_{t}+\eta\cdot(r_{\text{viol}}-r_{\text{target}})$ (24)
where $\eta>0$ is a step size and $r_{\text{target}}\in(0,1)$ is the desired maximum violation influence ratio. This adaptation increases $\rho$ when violation influence exceeds the target and decreases it otherwise, maintaining constraint satisfaction while avoiding overly conservative behavior. Algorithm 2 presents the complete online procedure.
Algorithm 2 Online: Datamodel-Based MPPI
0: Trained predictor $h_{\phi^{*}}$, initial state $x_{0}$, nominal control $\bar{U}$, initial penalty $\rho$, pruning threshold $\tau$, target violation ratio $r_{\text{target}}$, adaptation step size $\eta$, temperature $\lambda$
1:for each control iteration do
2:  Sample $K$ trajectories from current state
3:  Compute cost components: $(C_{\text{goal},k},C_{\text{ctrl},k},C_{\text{viol},k})$ for all $k$
4:  Compute total costs: $C_{k}\leftarrow C_{\text{goal},k}+C_{\text{ctrl},k}+\rho\cdot C_{\text{viol},k}$
5:  Compute instance statistics: $\bar{C}\leftarrow\frac{1}{K}\sum_{k}C_{k}$,  $\sigma_{C}\leftarrow\text{std}(\{C_{k}\})$
6:  Predict influence: $\hat{\theta}_{k}\leftarrow h_{\phi^{*}}(C_{k},C_{\text{viol},k},\bar{C},\sigma_{C})$ for all $k$
7:  Identify kept samples: $\mathcal{K}_{\text{keep}}\leftarrow\{k:|\hat{\theta}_{k}|\geq\tau\}$
8:  Identify violating samples: $\mathcal{K}_{\text{viol}}\leftarrow\{k:C_{\text{viol},k}>0\}$
9:  Compute violation ratio: $r_{\text{viol}}\leftarrow\frac{\sum_{k\in\mathcal{K}_{\text{viol}}}|\hat{\theta}_{k}|}{\sum_{k=1}^{K}|\hat{\theta}_{k}|}$
10:  Update penalty: $\rho\leftarrow\rho+\eta\cdot(r_{\text{viol}}-r_{\text{target}})$
11:  Compute weights: $\tilde{w}_{k}\leftarrow\frac{\exp(-C_{k}/\lambda)}{\sum_{j\in\mathcal{K}_{\text{keep}}}\exp(-C_{j}/\lambda)}$ for $k\in\mathcal{K}_{\text{keep}}$
12:  Compute optimal control: $u^{*}\leftarrow\sum_{k\in\mathcal{K}_{\text{keep}}}\tilde{w}_{k}\,u_{k}$
13:  Execute first control $u_{0}^{*}$, shift nominal sequence $\bar{U}$
14:end for
Remark 1 (Unified Influence for Efficiency and Safety).
Because the predictor takes both total cost and violation cost as input, a single learned model captures influence relevant to both sample efficiency and constraint handling. The pruning mechanism improves efficiency by discarding low-influence samples regardless of their constraint status, while the adaptive $\rho$ mechanism improves safety by strengthening penalties when constraint-violating samples remain influential. These two mechanisms operate complementarily: pruning reduces computation without sacrificing control quality, while penalty adaptation ensures the kept samples lead to safe behaviors.
Remark 2 (Computational Complexity).
The offline phase requires $O(N\cdot M\cdot K)$ subset evaluations and $N$ LASSO solves, but is performed only once. The online phase requires only $O(K)$ forward passes through $h_{\phi^{*}}$ per control iteration, making influence prediction negligible compared to trajectory sampling and cost evaluation. The pruning step further reduces the cost of the weighted average computation from $O(K)$ to $O(|\mathcal{K}_{\text{keep}}|)$.
IV Experiments
We evaluate the proposed framework on a path tracking task with obstacle avoidance, demonstrating unified benefits for sample efficiency and constraint-aware control.
IV-A Setup
We consider a vehicle governed by the kinematic bicycle model with state $x=[p_{x},p_{y},\psi,v]^{\top}$ (position, heading, velocity) and control $u=[\delta,a]^{\top}$ (steering angle, acceleration). The vehicle tracks an oval reference path while avoiding circular obstacles. The goal cost penalizes deviation from the reference path, while the violation cost applies a quadratic penalty when the vehicle enters the obstacle regions.
The MPPI controller uses a planning horizon of $T=20$ steps at $dt=0.1$ s and temperature $\lambda=100$. Control noise is sampled from $\mathcal{N}(0,\Sigma)$ with $\Sigma=\text{diag}(0.1,0.5)$ for steering and acceleration respectively.
For the influence predictor $h_{\phi}$, we use a multilayer perceptron (MLP) with two hidden layers of 64 units each and ReLU activations. The network takes as input the normalized costs $(S_{k},S_{\text{viol},k},\bar{S},\sigma_{S})$ and outputs the predicted influence coefficient $\hat{\theta}_{k}$. Training uses the Adam optimizer with learning rate $10^{-3}$ for 1000 epochs. The datamodel parameters are set as: inclusion probability $\alpha=0.5$, number of subsets $M=50$, and LASSO regularization $\mu=0.01$. We collect $N=200$ MPPI instances from randomly sampled initial states for offline training.
IV-B Sample Efficiency Analysis
Figure 2: Sample efficiency comparison between DM-MPPI and standard MPPI across varying sample sizes $K\in\{50,100,150,\ldots,500\}$.
To determine the minimum number of samples required for effective control, we evaluate both methods across sample sizes $K\in\{50,100,150,\ldots,500\}$. Figure 2 presents the results. Standard MPPI shows a steep performance degradation below $K=200$, with position RMSE increasing from 0.41 m at $K=500$ to 1.30 m at $K=50$. In contrast, DM-MPPI maintains stable performance across a wider range: at $K=100$, it achieves position RMSE of 0.44 m and $r_{\text{viol}}=0.11$, comparable to standard MPPI at $K=500$ (0.41 m and 0.09 respectively).
However, the datamodel approach has a minimum sample requirement. At $K=50$, DM-MPPI performance degrades noticeably (position RMSE increases to 0.72 m, $r_{\text{viol}}$ rises to 0.35), suggesting insufficient samples for reliable influence estimation. This aligns with the effective influence size $K_{\text{eff}}^{\theta}\approx 28$ observed in our analysis—when the total sample count approaches this threshold, the datamodel cannot effectively distinguish high-influence samples from noise.
Based on this analysis, we select $K=100$ for DM-MPPI in subsequent experiments, achieving $5\times$ sample reduction compared to standard MPPI ($K=500$) while maintaining equivalent control quality.
IV-C Performance Comparison
Figure 3: Comparison of sampled different trajectories.
TABLE I: Performance Comparison
Metric Standard DM-Fixed DM-Adaptive
Samples used 500 100 100
Position RMSE (m) 0.41 0.44 0.43
Heading RMSE (rad) 0.031 0.033 0.032
Min. obstacle dist. (m) 1.24 1.31 1.42
$r_{\text{viol}}$ 0.091 0.082 0.071
Iteration time (ms) 85.4 18.2 18.5
We compare three configurations: (1) Standard MPPI with $K=500$ samples and fixed penalty $\rho=10^{10}$, (2) DM-MPPI with $K=100$ samples and fixed $\rho=10^{10}$, and (3) DM-MPPI with $K=100$ samples and adaptive $\rho$ using $r_{\text{target}}=0.05$ and step size $\eta=10^{9}$.
Table I summarizes the performance comparison. The datamodel approach enables $5\times$ sample reduction while maintaining tracking quality. Figure 3 visualizes the sampled trajectories at a representative timestep; despite using only 100 samples, the datamodel variant produces a control output comparable to standard MPPI with 500 samples. Analysis of the learned influence coefficients reveals an effective influence size of $K_{\text{eff}}^{\theta}\approx 28$, confirming that roughly one-quarter of the samples carry meaningful influence while the remainder can be safely pruned. The adaptive variant achieves the largest minimum obstacle distance by dynamically adjusting $\rho$ based on the violation influence ratio $r_{\text{viol}}$.
V Limitations and Future Work
Soft Constraint Guarantees. The proposed framework relies on soft constraint penalties, which cannot guarantee hard constraint satisfaction. While the adaptive $\rho$ mechanism reduces the influence of constraint-violating samples, it cannot ensure that the resulting control will never violate constraints. Safety-critical applications may require integration with formal methods such as Control Barrier Functions [20, 21] or tube-based robust MPC [17]. Future work could integrate the datamodel framework with CBFs, using predicted influence to inform barrier parameters.
Limited Experimental Scope. The current evaluation considers a single path tracking task with obstacle avoidance using a kinematic bicycle model. While this serves as a representative benchmark, comprehensive validation requires evaluation across diverse domains including manipulation, legged locomotion, and higher-dimensional systems with more complex constraint structures. Future work will include comparisons against state-of-the-art sample-efficient MPPI variants [7, 16], evaluation under varying noise levels and constraint complexities, and deployment on real hardware platforms.
Fixed Predictor After Offline Training. The influence predictor is fixed after offline training, which limits adaptability to changing conditions. Online fine-tuning would allow the predictor to adapt to distribution shifts or novel scenarios, but introduces computational overhead that conflicts with real-time requirements. One promising direction is to maintain a small replay buffer of recent MPPI instances and perform incremental updates during idle cycles. Alternatively, meta-learning approaches could train the predictor to rapidly adapt to new conditions with minimal fine-tuning samples.
Generalization Across Cost Scales. The framework uses a fixed inclusion probability $\alpha$ during offline training, and the learned predictor generalizes well within the range of cost distributions encountered during training. However, performance may degrade when deployed in scenarios where cost statistics differ substantially from training. Normalizing costs by instance-level statistics partially addresses this issue, but does not fully resolve it when the underlying cost structure changes significantly. Future work could explore scale-invariant predictor architectures that operate on relative cost rankings rather than absolute values to enable broader generalization.
VI Conclusion
We presented a framework that extends the datamodeling paradigm from machine learning to sampling-based optimal control. While traditional datamodels estimate sample influence through regression on a fixed dataset, our approach trains an influence predictor that generalizes across control instances—enabling real-time influence estimation without solving regression problems online. By learning to map cost features to influence coefficients, the predictor identifies which samples contribute meaningfully to the control output, allowing efficient sample pruning and adaptive constraint handling through a single learned model.
The key insight is that influence coefficients, though computed via subset regression during offline training, show predictable structure as a function of sample costs and global statistics. This structure enables amortized inference: rather than re-running regression for each new MPPI instance, we predict influence directly from cost features in a single forward pass. Experiments on path tracking with obstacle avoidance demonstrate $5\times$ sample reduction while maintaining tracking performance and achieving improved safety margins through adaptive penalty tuning.